Astronomy Report – February 19th

Astronomy Report

Name: Eishi Kim
Crew: 222
Date: 2/19/2020


Robotic Telescope Requested: MDRS-WF (then MDRS-14 if approved)

Objects to be Imaged this Evening: M101 and M42 will be imaged as part
of the astrophotography project. In this regard, I’ve requested more
time on my Skynet account. If reverting back to the MDRS-14 is approved,
I will keep on with the still active observations from the previous sols.

Images submitted with this report: None

Problems Encountered: Imaging of M31 is to be abandoned for tonight, as
it won’t be visible long enough.


Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

Operations Report – February 19th

Crew 222 Operations Report 02/19/2020

SOL: 3

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Tuesday generator:
Off at 6:00am
On at 9:30pm SOC was 82%

Wednesday generator:
Off at 6:30am
On at 9:00pm SOC was XX%

Diesel Reading – 75%

Station Propane Reading – 37% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 47 gallons

Water Meter: 149334.72 units

Water (static tank) – 410 gallons

Static to Loft Pump used – yes

Water in GreenHab – 117.5 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes
Hours: 144.1
Beginning charge: 100%
Ending charge: 88%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: yes
Hours: 132.0
Beginning charge: 100%
Ending charge: 80%
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: no
Hours: 138.0
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery.
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 10 minutes
Notes on ATVs: The second one from the Hab doesn’t have anymore battery.

HabCar used and why, where? The habcar broke down on 02/18/2020. It’s
parked at Bull Mountain Market.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: used for the com. All ok.

Summary of suits and radios: Suits 7, 8, 9, 2 used and radio tested for
the EVA. The spacesuit number 10 was unplugged this morning so we didn’t
use it. All ok now.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: Operations on the spacesuit and LCVG experiment.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

GreenHab Report – February 19th

Crew 222 GreenHab Report – 19/02/2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Ambient with door open

Average temperatures: 22,6°C

Hours of supplemental light: NA

Daily water usage for crops: 9,5 gallons

Daily water usage for research and/or other purposes: 0 Liters

Water in blue Tank: 117,5 gallons

Time(s) of watering crops: 2

Changes to crops: NA

Narrative: Today, Since I was on EVA this morning, this is Marie, our
journalist that watered plants as I showed her yesterday and maked the
morning checks. After the lunch,
I started my experience of growing plants in gelatin to compare with
our culture tower. I went finally in the evening to watered a second
time the plants and took some
mint to prepare a tea to my crew this evening.

Harvest: 2g pineaple mint, 2g mint

Support/supplies needed: nothing to report

Sol Summary – February 19th

Crew 222 Sol Summary Report 19-02-2020
Sol: 3
Summary Title: Crew Spirit
Author’s name: Benjamin Auzou
Mission Status: Crewmembers taking marks in the station and with each
other, science running well.
Sol Activity Summary: In the morning, the crew woke up with an other
matinal workout session. Then four crewmember went on the first
exploration EVA of the mission, the EVA went well and they discovered
some landscapes that were new for them.
The afternoon were dedicated to the evolution of scientific experiments
: the Spacesuit developped by crew 206 and now improved by our engineer
Valentin will be ready in a few days for an EVA, the vertical garden is
now ready to receive plants.
Before dinner, we shared a moment all together, discussing on forces
that , and tonight we’re planning to watch a movie together. The crew
dynamic has evolved today, thanks to discussions together and the
moments shared in the Hab.
Look Ahead Plan:
Tomorrow, the crewmembers that stayed in the Hab this morning will go on
EVA to explore the south of the Hab.
Thanks to the monitor, we will be able to launch the water monitoring

Anomalies in work: none
Weather: blue sky in the morning, some clouds in the afternoon
Crew Physical Status: some weaknesses or physically tired
EVA: First exploration EVA
Reports to be filed: Journalist report, Operations Report, GreenHab
Report, Astronomy Report, EVA Report, EVA Request, Photo Report, HSO
Report, Photo of the Day
Support Requested: nothing to report

Journalist Report – February 19th

Crew 222 Journalist Report 19Feb2020

Author: Marie Bochard, Crew Journalist

Sol 3

Title: Explorers

Mars is truly a beautiful planet. This morning was our first exploration-focused EVA. The group left for North Ridge, and came across stunning sights. The red sand meets yellow stones, dark hills are intercut by lighter sediments, it seems every step brings a new landscape to our eyes. It feels like we are so far from Earth, and we are! Even with spacesuits on, we all wished to have explored the surroundings of the Hab for hours on end.

Today we take a tour of the GreenHab with Bleuenn, our GreenHab officer. She is responsible for watering, planting and harvesting all the plants that we have. The GreenHab is our own little piece of Earth, a reminder of the green and blue planet that we left. Having this greenery is extremely important to us. Everything around us is hostile and infertile, whereas home grows in the soil of the GreenHab. It plays a big part in cheering us up: seeing plants grow and eating fresh food is very good for the morale. Back home, we are all used to eating fresh fruits and vegetables at every meal, especially since some of us are vegetarians, and the dehydrated food we have doesn’t have the same texture as food back on Earth. Flowers can also help with the isolation, as they can bring the smell of home to the Hab. All in all, the GreenHab is not just a habitat for vegetation, it also helps keep everyone in good mental shape!

We ended the day with a group cohesion activity. Nicolas spread cards on the table with qualities on them, and we each chose a couple that represented us. The goal was to explain how these qualities defined us, using examples from our experience. This was a very enriching experience as we got to know more of everyone, and being able to open up to the crew is important to be fully integrated. It was a great way to end the day, we are all feeling good and ready for tomorrow

EVA Report – February 19th

Crew 222 EVA Report 19-02-2020
EVA # 3
Author: Bleuenn ROIRON (GHO)
Purpose of EVA: Change batteries of LOAC and MegaAres, and then first
exploration for the crew 222
Start time: 9:04 AM
End time: 11:07 AM
Narrative: First, we went out to check all the rovers and the water. We
moved then in teams of two to our experience site to change
batteries and check the data for the MegaAres. Everything went fine so
we moved to URC Northsite and arrived at 9:45 to start our exploration.
All of us were very excited for that. We walked for approximately one
hour bewteen URC Northsite and North Ridge. Landscapes were really
differents from those we are used to, especially because of colors, the
ground was red, beige and sometimes almost purple. We climbed a
small hill to have a large overview on what was around us and took a
break on it to really appreciate landscapes. Then we came back to the hab.
Before coming back in it, we had a small check on tunnels. We entered
the station again at 11:07 AM.
Destination: URC Northsite
Coordinates (use UTM NAD27 CONUS):4251800/519000
Participants: Bleuenn ROIRON(GHO), Nina SEDBON(crew Scientist), Nicolas
BOURDEAUD(HSO), Valentin PRUDHOMME(crew engineer)
Road(s) and routes per MDRS Map: Cow Dung Road
Mode of travel: Rovers : Spirit & Deimos

Copyright © The Mars Society. All rights reserved. | Main Site